Space Stations

NASA ISS On-Orbit Status 4 June 2014

By Marc Boucher
Status Report
June 5, 2014
Filed under , ,
NASA ISS On-Orbit Status 4 June 2014
NASA ISS On-Orbit Status 4 June 2014.

Today: Advanced Colloids Experiment (ACE)-M-2 Sample Start: Flight Engineer (FE)-5 Wiseman configured the Light Microscopy Module (LMM) in the Fluids Integrated Rack (FIR) and inserted the second sample for an ACE-M-2 investigation run.
In the ACE-M-2 investigation, a model colloidal system is used to observe the microscopic behavior of liquids and gases separating from each other, near the critical point, via a process called spinodal decomposition.

Ocular Health: Wiseman assisted Commander (CDR) Swanson and FE-6 Gerst in taking ultrasound images of their eyes. All three USOS crewmembers took cardiac echo measurements of each other as part of the Ocular Health data collection. The ultrasound images will be used to identify changes in globe morphology, including flattening of the posterior globe, and document optic nerve sheath diameter, optic nerve sheath tortuosity, globe axial measurements, and choroidal engorgement. The Ocular Health protocol calls for a systematic gathering of physiological data to characterize the risk of microgravity-induced visual impairment/intracranial pressure in ISS crewmembers. Researchers believe that the measurement of visual, vascular and central nervous system changes over the course of this experiment and during the subsequent post-flight recovery will assist in the development of countermeasures, clinical monitoring strategies, and clinical practice guidelines.

Microbiome: Gerst collected blood and saliva samples for his Flight Day (FD) 7 Microbiome session. The Microbiome experiment investigates the impact of space travel on both the human immune system and an individual’s microbiome (the collection of microbes that live in and on the human body at any given time).

EVA Mobility unit (EMU) 3010 Return to Service: EMU 3010 is being returned to service via a series of activities this week. Today Swanson installed a cable to both the EMU radio and antenna. Tomorrow he will perform a series of checkouts to determine if the EMU is ready for use.

Hardware Control Panel (HCP) Setup: As part of Launch Commit Criteria (LCC), today Gerst set up the HCP in the Japanese Pressurized Module (JPM) in preparation for Cygnus Orbital-2 arrival planned for later this month. The HCP will be used if a manual abort is required during Cygnus rendezvous.

Pumping from USOS Urine Containers [???-?] to 53Progress (53P) Tanks: Today, the Russian crew was supposed to pump 6 ???-?s into the 53P Rodnik tanks. They were only able to pump 2 ???-?s. The Russian team is evaluating why the additonal ???-?s did not fit.

Today’s Planned Tasks

RSS1,2 reboot / r/g 1009 [Complete]
Laptop RS1(2) reboot [Complete]
CHROMATOMASS. Science Ops prep / r/g 5791 [Complete]
Biochemical urine test [Complete]
CHROMATOMASS. Saliva and blood sampling / r/g 5791 [Complete]
HRF – sampling and stowage [Complete]
Biochemical urine test [Complete]
HRF – stow samples into MELFI [Complete]
HRF – centrifuge activation and setup [Complete]
URISIS stow [Complete]
??? SM Testing [Complete]
HRF- blood draw assist[Complete]
HRF – cooler centrifuge configuration [Complete]
DPC (S-band) [Complete]
Preliminary EVA-38 timeline review,r/g 5792 [Complete]
HRF – retrieve blood samples and deactivate cooler centrifuge [Complete]
HRF samples installation in MELFI [Complete]
HRF – h/w ops closeout ops and stowage [Complete]
FMK atmospheric formaldehyde monitor de-install[Complete]
FIR – rack doors opening [Complete]
CHROMATOMASS. Closeout ops / r/g 5791 [Complete]
LMM – microscope setup [Complete]
CHROMATOMASS SPECTRE M. r/g 5791 [Complete]
ACE – sample installation [Complete]
Daily arterial pressure measurement (completion) r/g 5785 [Complete]
BIOME – questionnaire fill-out [Complete]
VIRTUAL. Experiment run. r/g 5790 [Complete]
CCD-camera setup Resist Tubule [Complete]
LMM – microscope setup [Complete]
FIR – rack doors closing [Complete]
CMS kits inventory / r/g 5793 [Complete]
Treadmill fitness evaluation / r/g 5799 [Complete]
WHC – fill-out [Complete]
CCAA cycling [Complete]
Treadmill fitness evaluation / r/g 5799 [Complete]
Set up h/w and connect HCP control panel power and command cables [Complete]
USND2 – h/w power ON [Complete]
Ultrasound scanning – h/w setup [Complete]
Ultrasound scanning – assist [Complete]
Ultrasound scanning – subject [Complete]
Cygnus – PROX and ??? power ON [Complete]
Ultrasound scanning – data transfer [Complete]
Use CMS to adjust [????-4?] CO ratio. r/g 5794 [Complete]
On MCC Go Urine transfer from USOS to Progress 421 Rodnik ??2 tank / r/g 5795 [Complete]
Familiarization with ACS (auxiliary computer systems). r/g 5796 [Complete]
??-1 sensor setting c/o [Complete]
Ultrasound scanning – subject [Complete]
Ultrasound scanning – assist [Complete]
EMU antenna checkout, inspection and photography [Complete]
??? maintenance [Complete]
OTKLIK. H/w activation.r/g 5797 [Complete]
Ultrasound scanning – data transfer [Complete]
Transfer items for disposal to Progress #421(??), IMS update / r/g 5685, 5750 [Complete]
Ultrasound scanning – closeout ops [Complete]
IMS update [Complete]
SHD – daily headache monitoring questionnaire [Complete]
Crew time for adaptation and familiarization with the station [Complete]
WRS water sample analysis [Complete]
SHD – daily headache monitoring questionnaire [Complete]
Photo/TV hardware questionnaire [Complete]
OH – h/w setup [Complete]
Crew time for adaptation and familiarization with the station [Complete]
OH – blood pressure measurement assistance [Complete]
Urine transfer from USOS to Progress 421 Rodnik ??2 tank / r/g 5795 [Partially Complete]
BIOME – sampling prep [Complete]
VEG-01 – water fill for plant watering [Complete]
OH – setup [Complete]
OH – blood pressure measurement assistance [Complete]
“Hematocrit”. Hardware prep [Complete]
DPC (S-band) [Complete]
TOCA – data recording [Complete]
SHD – daily headache monitoring questionnaire [Complete]
USND2 – power OFF [Complete]
JRNL- journal entry [Complete]
Task List

ARED Cylinder Evac
IFM RPCM TL/jumper gather
WHC KTO replace
Roscosmos Report
Uragan photo
BASS ops overview
BASS single TL ops parts 1,2,3
CCE CDR [Complete]
CSL Client name-verify [In Work]
BCATC1 sample 8 pre-mix
CEVIS Isolator MOD
Cygnus pre-gather
Crew handover
RS EVA H/W gather
CFE ICF5 ops parts 1,2,3
EMU LLB Auto-Init
EMU LLB SSC Relocate
EVA ORU Bag Inspection
IFM MOOT restow
NOD2O2 Cleanout
JPM HD assessment
NOD3 PEPS assessment
ODF CTB reconfig
P/TV H/W questions part 2

Nominal commanding
Three-Day Look Ahead:

Thursday, 06/05: EMU 3010 Return To Service, Prox HCP checkout, LA1B RPCM R&R
Friday, 06/06: HRF blood collection, OBT ISS Emergency h/w familiarization, OBT Cygnus rendezvous, Cygnus pre-gather
Saturday, 06/07: Crew off duty
QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – On
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – Off
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Override
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Stop
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Standby
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.